Statsitics 36-303: Sampling, Surveys and Society Tentative Schedule of Topics

• Following is a tentative schedule of topics for the course.

Please check http://www.stat.cmu.edu/˜ brian/303 regularly, for the latest information.

• Especially after spring break, the schedule may change somewhat, due to Dr. Junker’s travel, and/or guest speakers and other events that may be scheduled later.

• Please note the dates marked for in-class presentations (**). Each group will present a progress report in Weeks 10–11, and a final report in weeks 14–15.

Week  Date  Tentative Topics  Tentative Sources
Week 1 Jan 11, 13  Introduction, Projects  Class notes, handouts, Groves Ch

1

Week 2 Jan 18, 20 Components of a Survey, Survey Quality Groves Ch’s 2, 3 Week 3 Jan 25, 27 Sampling Methods, Survey Ethics

G&H Ch’s 5, 11

Week 4 Feb 01, 03^{∗}  Statistics of Surveys  Handouts, Groves Ch 4
Week 5 Feb 08, 10  Survey Questions  Groves, Ch’s 7, 8

Week 6 Feb 15, 17 Review (2/15), MIDTERM EXAM (2/17) Week 7 Feb 22, 24 The Analysis of Stratified Samples Handouts, Groves Ch 4 Week 8 Mar 01, 03 Dealing with Nonresponse & Missing Data

Groves, Ch 6

Week 8 Mar 08, 10 SPRING BREAK

HAVE FUN!

Week 9 Mar 15, 17 The Analysis of Clustered Samples Handouts, Groves Ch 4 Week 10 Mar 22^{∗∗}, 24^{∗∗} Project Progress Reports

Groves, Ch 12

Week 11 Mar 29^{∗∗}, 31 Post-Survey Processing

Groves, Ch 10

APRIL 1 SUBMIT ROUGH DRAFT OF PROJECT

REPORT

APRIL 1 DEADLINE FOR MEETING OF THE MINDS

http://www.cmu.edu/uro/MoM

Week 12 Apr 05, 07 Topics/Catch-up; Review Week 13 Apr 12, 14 FINAL EXAM (4/12); Carnival (4/14)

TBA

Week 14 Apr 19, 21^{∗∗} Topics TBA

Week 15 Apr 26^{∗∗}, 28^{∗∗} Project Final Reports

TBA

APRIL 29 SUBMIT FINAL DRAFT OF PROJECT

REPORT

MAY 4 MEETING OF THE MINDS http://www.cmu.edu/uro/MoM

∗ Dr. Junker absent this day. Stay tuned for other arrangements. ∗∗ In-class presentation days. There will be three presentations per lecture.